OpenStudy (anonymous):

x − y = 10 x + y = 6 (−2, 8) (−8, 2) (8, −2) (8, 2)

jimthompson5910 (jim_thompson5910):

Hint: add the equations to get x-y = 10 x+y = 6 --------- 2x+0y = 16 so 2x = 16

OpenStudy (anonymous):

so if you rearrange you see x=10+y x+6-y so 10+y=6-y solve for y and sub back into the equation to find x :)

OpenStudy (anonymous):

sorry @jim_thompson5910 my computer doesn't seem to show me when other people respond lol but yes @Marcus980 jim is right!!

jimthompson5910 (jim_thompson5910):

that's ok sarahusher

OpenStudy (anonymous):

When I slove For Y I Got -2

jimthompson5910 (jim_thompson5910):

y = -2 is correct

OpenStudy (anonymous):

(−2, 8) (−8, 2) (8, −2) (8, 2)

OpenStudy (anonymous):

remember coordinates are in the form (x,y) so just sub in your values :)

OpenStudy (anonymous):

So (-2,8)

OpenStudy (anonymous):

double check which comes first, x then y if x=8 y=-2 (x,y)

OpenStudy (anonymous):

I sse What You Are Talking About It The 8,-2

OpenStudy (anonymous):

Yes! Good work